OpenStudy (anonymous):

mike wants to make a meatloaf. his recipe uses a total of 3 pounds of meat if he use a 3 to 1 ratio of beef pork, how much pork will he use? enter the answer as a mixed number in the simplest form

OpenStudy (phi):

I assume you mean 3 to 1 ratio of beef to pork they tell us he uses total of 5 pounds of meat (which means the sum of the beef+pork is 5) to answer how much pork ? we need a ratio of pork to total (not beef to pork) we get total by adding up the 3 and 1 in 3 to 1 ratio we get 4 write the ratio as a fraction \[ \frac{1}{4} = \frac{pork}{total} \] fill in 5 for total, and solve for pork \[ \frac{1}{4} = \frac{pork}{5} \]

OpenStudy (anonymous):

@phi this the same question but different pounds